OpenStudy (anonymous):

Simplify the following expression. sqrt-104 sqrt -8

OpenStudy (anonymous):



OpenStudy (mertsj):

\[\sqrt{-104}\sqrt{-8}=\sqrt{4(2)(13)(4)(2)i^2}=-8\sqrt{13}\]

OpenStudy (anonymous):

Thanks I see where I made my mistake at.

OpenStudy (mertsj):

yw